# ASSESS: A Semantic and Structural Evaluation Framework for Statement Similarity

### v1

@@ 3.3.2 Implementation
Since any practical implementation is limited to a finite list of transformations, our algorithm computes a tractable upper bound of the TransTED, rather than its exact theoretical value. However, if this finite sequence of transformations proves that two expressions are semantically equivalent, the computed distance is exactly 0 - the true value, not an approximation.
@@
apply forall_congr;
intro _
(a $\rightarrow$ b) = (a $\rightarrow$ c) (_ : a) $\vdash$ b = c
apply implies_congr;
all_goals (try exact?)  [^1] At least one of a = b and c = d should be proved by exact?, or otherwise the tactic command generates two new goals, which is forbidden in our heuristic search.
(a $\rightarrow$ c) = (b $\rightarrow$ d) a = b or c = d or None
ext
(fun (x : A) => f x) =
(fun (y : A) => g y)
(x : A) $\vdash$ f x = g x
@@
In the context of the Lean theorem prover, transformation can be conceptualized as the set of all tactics that adhere to our theoretical framework (see Theorem 1). For implementation, on the one hand, we curated a subset of these tactics, selected for their practical efficacy and detailed in Table 1. On the other hand, since these tactics operate on single equalities rather than pairs of statements, we first construct a unified expression for comparison. Given a pair of formal statements, we achieve this by forming an equality between their respective types. This process converts the statement pair into a single equation, which then serves as the input for our transformation tactics. Some concrete examples are provided in Appendix C.2.
To circumvent the combinatorial explosion of applying all tactics, we apply a search algorithm. The algorithm performs a heuristic search through transformations. The guiding heuristic prioritizes tactics that reduce the TED between OPTs on the left- and right-hand sides of the equation, effectively pruning less promising branches. This approach allows our method to efficiently find short-distance transformation paths and the complete algorithm is detailed in Appendix B. Finally, to enhance interpretability, we normalize the raw TransTED distance into a TransTED Similarity score, defined in an analogous way as Definition 3.
@@ 4.1 The EPLA Benchmark
We introduce EPLA (Evaluating Provability and Likeness for Autoformalization), a new benchmark designed to overcome the limitations of current autoformalization evaluations. Existing benchmarks often rely on coarse binary labels (e.g., True/False), which prevent a nuanced assessment of a metric’s performance. For instance, a slightly incorrect but easily fixable translation is more valuable than one that is completely wrong, yet both would be identically classified as False. EPLA provides a more granular, versatile, and interpretable evaluation to address this critical gap.
EPLA is built upon two established datasets: miniF2F-test (Zheng et al., 2022) and ProofNet-test (Azerbayev et al., 2023). We utilize the specific versions of these datasets provided by Numina [^3] https://huggingface.co/datasets/AI-MO/minif2f_test for miniF2F-test and DeepSeek [^4] https://github.com/deepseek-ai/DeepSeek-Prover-V1.5/tree/main/datasets for ProofNet-test. To generate candidate formalizations, we translate the natural language statements from these source datasets into Lean 4 using two distinct translators: the Herald Translator (Gao et al., 2024) and Gemini 2.5 Pro (Comanici et al., 2025). After applying a compilation filter to retain only syntactically valid statements, a panel of five experts with backgrounds in mathematics annotated and cross-validate each formal statement pair. The annotation was performed along two axes: Semantic Provability, defined as whether the ground-truth and predicted statements are mutually provable, and Structural Likeness, which measures the degree of structural likeness between them. The final EPLA benchmark, detailed in Table 2, comprises 524 such annotated data points.
@@
Herald Translator Gemini 2.5 Pro Total
EPLA-miniF2F 198 175 373
EPLA-ProofNet 89 62 151
Note: The intersection of these axes yields EPLA’s five-label annotation scheme: (a) Mutually Provable & Like; (b) Mutually Provable & Unlike; (c) Mutually Unprovable & Like; (d) Mutually Unprovable & Like* (where Like* denotes likeness after semantics-preserving transformations); and (e) Mutually Unprovable & Unlike. This granular classification enables flexible evaluation regimes. For instance, a strict evaluation could define success as categories (a) and (b), whereas a human-in-the-loop scenario might accept (a) through (d) as useful outputs.
@@ 4.2 Experiment Setting
  • Typecheck: A predicted formal statement is deemed correct if it successfully compiles.
@@
  • Majority Voting: Following Lean Workbook (Ying et al., 2024a), we employ DeepSeek-V3.1 (Liu et al., 2024) with temperature $0.7$ for 16 rounds of majority voting.
  • Definitional Equality (Liu et al., 2025a): A predicted formal statement is considered correct under this metric if example: thm_P = thm_Q := by rfl succeeds.
@@
To ensure a fair comparison against baselines that require binary labels, we established a unified evaluation protocol. First, we map EPLA’s granular annotations to a binary ground truth, designating categories (a) and (b) as True and all others as False. Second, for continuous metrics (BLEU, TED Similarity, and TransTED Similarity), we convert their scores to binary predictions by applying a threshold and the best results over all possible thresholds are reported.
All experiments were conducted on our EPLA benchmark using the Lean toolchain v4.9.0-rc1. With the sole exception of the GPU-dependent BEq baseline, all evaluated metrics were executed on CPUs. Our proposed metric is particularly lightweight, requiring only interaction with the Lean Language Server and REPL, thus imposing lower computational overhead.
Table 3: Overall results of the competing baselines and our metrics. The boldface refers to the highest score and the underline indicates the next best result of the metrics. Detailed results about the number of TP, TN, FP and FN are available in Appendix C.1.
@@
Identity Match 100.00% 9.78% 33.24% 0.05 100.00% 1.25% 47.68% 0.01
Typecheck 73.99% 100.00% 73.99% 0.00 52.98% 100.00% 52.98% 0.00
BLEU 89.69% 63.04% 67.29% 0.33 64.21% 76.25% 64.90% 0.29
Majority Voting 89.95% 61.59% 66.49% 0.33 72.86% 63.75% 68.21% 0.37
Definitional Equality 98.92% 33.33% 50.40% 0.20 80.00% 10.00% 50.99% 0.07
BEq 100.00% 48.91% 62.20% 0.33 100.00% 28.75% 62.25% 0.28
TED Similarity 85.83% 74.64% 72.12% 0.35 65.98% 80.00% 67.55% 0.34
TransTED Similarity 86.08% 85.14% 78.82% 0.46 67.31% 87.50% 70.86% 0.40
@@ 4.3 Experiment Results
Table 3 presents a comprehensive overview of our proposed TransTED Similarity compared to various baseline metrics across the EPLA-miniF2F and EPLA-ProofNet benchmarks. Our primary evaluation focuses on accuracy and the Cohen’s Kappa, as accuracy directly reflects the overall correctness of the evaluation, and Kappa (Cohen, 1960) offers a robust measure of agreement beyond chance. The results demonstrate that TransTED Similarity achieves state-of-the-art performance across both datasets. On EPLA-miniF2F, it secures the highest accuracy (78.82%) and Kappa (0.46). It maintains this superior performance on the EPLA-ProofNet dataset, again leading with the highest accuracy (70.86%) and Kappa (0.40). This consistent lead, particularly in the Kappa metric, highlights our method’s robust ability to align with ground-truth judgments, outperforming all evaluated baselines.
@@
As detailed in Table 3, TransTED Similarity overcomes key limitations observed across different categories of baseline methods. First, while proof-based metrics like Definitional Equality and BEq achieve high precision, their performance is hampered by low recall. This brittleness stems from the inherent limitations of current Automated Theorem Provers (ATPs), which often fail to prove valid equivalences and thus produce a high rate of false negatives. Our method avoids this rigid dependency, offering a much more balanced assessment. Furthermore, TransTED Similarity significantly outperforms the LLM-based Majority Voting baseline, achieving superior accuracy (78.82% vs. 66.49% on EPLA-miniF2F) and a substantially higher Kappa score (0.46 vs. 0.33). Crucially, our method also sidesteps the primary drawbacks of LLM-based metrics: it is computationally inexpensive and fully reproducible. Finally, our method addresses a key weakness of n-gram-based metrics like BLEU: sensitivity to threshold tuning. As illustrated in Figure 3, TransTED Similarity is remarkably robust, maintaining high and stable performance across a wide range of decision thresholds. This stability is critical for practical application, as it ensures reliable outcomes without the need for costly, dataset-specific threshold optimization.
Figure 3: Comparison of BLEU and TransTED Similarity across thresholds on EPLA-miniF2F and EPLA-ProofNet. The plots illustrate that TransTED Similarity consistently achieves a higher Kappa score than BLEU and exhibits far greater stability across a wide range of decision thresholds.
@@
We conduct an ablation study comparing our TransTED Similarity metric against its baseline, TED Similarity. The results, presented in Table 3, demonstrate that this component provides a significant performance boost on both datasets. On EPLA-miniF2F, its inclusion increases accuracy by +6.70 percentage points from 72.12% to 78.82% and raises the Kappa score from 0.35 to 0.46. This trend holds on EPLA-ProofNet, where accuracy improves by +3.31 percentage points (67.55% → 70.86%) and Kappa increases from 0.34 to 0.40. These consistent gains validate our central hypothesis: by integrating semantic transformations, our metric can successfully recognize statements that are semantically equivalent despite being syntactically diverse. This capability allows TransTED Similarity to achieve a more accurate and human-aligned evaluation where purely structural methods like the standard TED Similarity falter.
@@ 5 Conclusion
In this work, we addressed the critical gap in evaluating autoformalized statements, where existing metrics fail to adequately balance semantic and structural information. We introduced ASSESS, a novel framework whose core contribution is the TransTED Similarity metric. By augmenting the structural comparison of Operator Trees with a set of semantic transformations, our method provides a more holistic and robust measure of statement similarity. We also developed EPLA, the largest expert-annotated benchmark for this task to date. Experiments conducted on EPLA demonstrate that TransTED Similarity establishes a new state-of-the-art, outperforming existing methods in both accuracy and agreement with human judgment (Kappa). This work provides the autoformalization community with a reliable, efficient, and reproducible metric for automated evaluation.
Appendix A Proof of Theorem 1

### v2

@@ 3.3.2 Implementation
Since practical implementations are constrained by a finite set of transformations, our algorithm computes a tractable upper bound of TransTED rather than its exact theoretical value. Crucially, however, if the transformation sequence establishes semantic equivalence, the computed distance becomes exactly 0, coinciding with the true theoretical value.
@@
apply forall_congr; intro _ (a $\rightarrow$ b) = (a $\rightarrow$ c) (_ : a) $\vdash$ b = c
apply implies_congr; all_goals (try rfl)  [^1] At least one of a = b and c = d should be proved by rfl, or otherwise the tactic command generates two new goals, which is forbidden in our heuristic search. (a $\rightarrow$ c) = (b $\rightarrow$ d) a = b or c = d or None
ext (fun (x : A) => f x) = (fun (y : A) => g y) (x : A) $\vdash$ f x = g x
@@
In the context of the Lean theorem prover, transformation can be conceptualized as the set of tactics that adhere to our theoretical framework (see Theorem 1). For implementation, on one hand, we curated a set of tactic commands, including rw?, the most related automatic tactic for searching equivalence transformations, and additional ones selected for their practical efficacy, listed in Table 1. On the other hand, since these tactics operate on single equalities rather than pairs of statements, we first construct an equation as the initial contrast to be transformed and unified later, by connecting a given pair of formal statements with an equal sign. Concrete examples are provided in Appendix F.3.
To circumvent the combinatorial explosion of applying all tactics, we apply a search algorithm. The algorithm performs a heuristic search through transformations. The guiding heuristic prioritizes tactics that reduce the TED between OPTs on the left- and right-hand sides of the equation, effectively pruning less promising branches. Searching is terminated as long as one of the following conditions is satisfied: The equivalence is proved, the node limit is exceeded (NLE), and the time limit is exceeded (TLE). This approach allows our method to efficiently find short-distance transformation paths and the complete algorithm is detailed in Appendix D. Finally, we normalize the raw TransTED distance into a TransTED Similarity score, defined in an analogous way as Definition 3.
@@ 4.1 The EPLA Benchmark
We introduce EPLA (Evaluating Provability and Likeness for Autoformalization), a benchmark designed to transcend the limitations of current evaluation protocols. Existing benchmarks frequently rely on coarse binary labels (e.g., provable vs. unprovable), precluding a nuanced assessment of metric performance. For instance, a slightly incorrect yet easily fixable translation is significantly more valuable than a completely erroneous one, yet binary metrics classify both identically as False. EPLA provides a more nuanced evaluation standard to address this critical gap.
EPLA is built upon two established datasets: miniF2F-test (Zheng et al., 2022) and ProofNet-test (Azerbayev et al., 2023). We utilize the specific versions of these datasets provided by Numina [^3] https://huggingface.co/datasets/AI-MO/minif2f_test for miniF2F-test and DeepSeek [^4] https://github.com/deepseek-ai/DeepSeek-Prover-V1.5/tree/main/datasets for ProofNet-test. To generate candidate formalizations, we translate natural language statements from the datasets into Lean 4 using four distinct models: two domain-specific models, Herald Translator (Gao et al., 2025) and Goedel-Formalizer-V2-8B (Lin et al., 2025), and two general-purpose models, Gemini-2.5-Pro (Comanici et al., 2025) and Qwen3-Max (Yang et al., 2025). The prompts used for the domain-specific models are consistent with the original papers, while the prompts used for the general-purpose models are provided in Appendix F.5.
Following the application of a compilation filter to retain only syntactically valid statements, a panel of seven experts with mathematical backgrounds annotated and cross-validated each pair of formal statements. The annotation methodology employs three Boolean sub-labels: Provability, representing the semantic equivalence of the statements; Likeness before transformation, indicating structural similarity between the original statements; and Likeness after transformation, indicating structural similarity following semantic-preserving transformations. The resulting EPLA benchmark, detailed in Table 2, comprises 1,247 annotated instances. Further specifications regarding the benchmark’s format and label definitions are provided in Appendix E.
@@
Herald Translator Goedel-Formalizer Gemini 2.5 Pro Qwen3-Max Total
EPLA-miniF2F 198 234 175 224 831
EPLA-ProofNet 89 130 62 135 416
@@ 4.2 Experiment Setting
  • Majority Voting: Following Lean Workbook (Ying et al., 2024a), we employ DeepSeek-V3.2-Exp (DeepSeek-AI, 2025) with temperature $0.7$ for 16 rounds of majority voting.
  • Definitional Equality (Liu et al., 2025a): A predicted formal statement is considered correct if example: thm_P = thm_Q := by rfl succeeds.
@@
The prompts employed in the Majority Voting baseline are detailed in Appendix F.5. Regarding tactic usage for the BEq baseline, we adopt the normal set of tactics exact, exact?, have, apply, cases’, constructor, ext, intro, intros, rw, use and fix the number of LLM generation attempts at $k=16$, consistent with the optimal hyperparameters reported in the original work. For TransTED Similarity, we utilize the top-5 suggestions generated by the rw? tactic at each step of the search process and define the termination criteria as a maximum tree size of 32 or a search timeout of 10 minutes.
To ensure a fair comparison with baselines requiring binary labels, we establish a unified evaluation protocol. First, we utilize the provability labels from EPLA as the ground-truth labels for the formal statement pairs. Second, for real-valued metrics such as BLEU, TED Similarity, and TransTED Similarity, we convert scores into binary predictions via thresholding. We report the optimal Kappa score and associated statistics, determined by sweeping over all possible decision thresholds.
All experiments were conducted on the EPLA benchmark using the Lean toolchain v4.9.0-rc1. The computational environment consisted of one NVIDIA A800 GPU with 80GB of memory and two Intel Xeon Silver 4216 CPUs, providing a total of 32 physical cores. With the exception of the GPU-dependent BEq baseline, all evaluations were executed on CPUs. Our proposed metric is particularly lightweight, requiring only interaction with the Lean Language Server and REPL.
Table 3: Overall results of the competing baselines and our metrics. The boldface refers to the highest score and the underline indicates the next best result of the metrics. Detailed results about the number of TP, TN, FP and FN are available in Appendix F.2.
@@
Identity Match 100.00% 8.79% 32.61% 0.05 100.00% 3.69% 43.51% 0.03
BLEU 82.25% 73.94% 68.96% 0.26 72.30% 43.85% 57.21% 0.18
Majority Voting 91.00% 31.27% 46.93% 0.14 66.67% 45.08% 54.57% 0.12
Definitional Equality 97.09% 32.57% 49.46% 0.19 86.21% 10.25% 46.39% 0.07
BEq 98.60% 45.77% 59.45% 0.29 98.77% 32.79% 60.34% 0.28
TED Similarity 84.51% 71.99% 69.56% 0.31 66.11% 81.56% 64.67% 0.23
TransTED Similarity 87.97% 69.06% 70.16% 0.35 68.49% 81.97% 67.31% 0.30
@@ 4.3 Experiment Results
Table 3 presents a comprehensive overview of our proposed TransTED Similarity compared to various baseline metrics across the EPLA-miniF2F and EPLA-ProofNet benchmarks. Our primary evaluation focuses on accuracy and the Cohen’s Kappa, as accuracy directly reflects the overall correctness of the evaluation, and Kappa (Cohen, 1960) offers a robust measure of agreement beyond chance. The results demonstrate that TransTED Similarity achieves state-of-the-art performance across both datasets. On EPLA-miniF2F, it secures the highest accuracy (70.16%) and Kappa (0.35). It maintains this superior performance on the EPLA-ProofNet dataset, again leading with the highest accuracy (67.31%) and Kappa (0.30). This consistent lead, particularly in the Kappa metric, highlights our method’s robust ability to align with ground-truth judgments, outperforming all evaluated baselines.
@@
As detailed in Table 3, TransTED Similarity effectively mitigates key limitations inherent to existing baseline categories. First, while proof-based metrics like Definitional Equality and BEq achieve high precision, their performance is plagued by low recall. This brittleness stems from the limitations of current Automated Theorem Provers, which frequent fail to verify valid equivalences, resulting in a high rate of false negatives. Our method sidesteps this rigid dependency, yielding a significantly more balanced assessment (see Appendix B for a comprehensive discussion). Furthermore, TransTED Similarity significantly outperforms the Majority Voting baseline, achieving superior accuracy (70.16% vs. 46.93% on EPLA-miniF2F) and a substantially higher Kappa score (0.35 vs. 0.14). Crucially, our method circumvents the primary limitations of LLM-based metrics: it is GPU-independent and fully reproducible. Finally, our method overcomes a critical limitation of n-gram-based metrics like BLEU: their inability to capture the semantics of mathematical text. As illustrated in Figure 3, TransTED Similarity demonstrates significant robustness, maintaining high stability across a broad spectrum of decision thresholds. Further details regarding threshold selection are provided in Appendix F.1.
Figure 3: Comparison of BLEU, TED Similarity, and TransTED Similarity across varying decision thresholds on EPLA-miniF2F and EPLA-ProofNet. Red stars indicate the optimal threshold configuration for each metric. Details regarding threshold selection are provided in Appendix F.1.
@@
We conduct an ablation study comparing our TransTED Similarity metric against the basic TED Similarity metric. The results, presented in Table 3, demonstrate that this component provides a significant performance boost on both datasets. On EPLA-miniF2F, its inclusion increases accuracy by 0.60 percentage points, improving it from 69.56% to 70.16%, and raises the Kappa score from 0.31 to 0.35. This trend holds on EPLA-ProofNet, where accuracy improves by 2.64 percentage points from 64.67% to 67.31%, and the Kappa score increases from 0.23 to 0.30. These consistent gains substantiate our central hypothesis: by integrating semantic transformations, our metric effectively discerns semantic equivalence amidst syntactic diversity. This capability enables TransTED Similarity to deliver a more accurate and human-aligned evaluation, succeeding where purely structural methods, such as the standard TED Similarity, fall short.
Figure 4: Statistic of the usage frequencies and adoption rates of tactic commands. (A) Usage Percentage (total): The proportion of each tactic command’s successful applications out of all successful applications during the entire search process. (B) Usage Percentage (adopted): The proportion of each tactic command’s adoption counts out of all adoption counts. (C) Adoption Rate: The proportion of each tactic command’s adoption counts out of its total application counts.
Analysis of Tactic Commands.
To assess the practical utility of our tactic set, we conducted a detailed statistical analysis of their usage patterns during the search process. Specifically, we quantified the frequency with which each tactic command was adopted. Here, an adopted tactic is defined as one applied along the optimal transformation path that yields a minimum TED state. These statistics are summarized in Figure 4, with raw data in Appendix F.2.
High usage frequency of rw? and norm_cast indicates their utility as general-purpose tools for simplification through rewriting and type-coercion normalization, and their low adoption rate is consistent with their role in exploring a wide search space and handling routine transformations. Conversely, high adoption rates of apply forall_congr; intro _ and rw [propext and_imp] illstrate their high specificity: they are applied less often, but when their preconditions - decomposing a universal quantifier or restructuring a nested implication - are met, they are highly effective and frequently reside on the critical path towards simplification. Other tactics commands are rarely successfully used, due to their strict application requirements and strong dependence on precise type unification, limiting their general applicability.
Overall, the synergy between high-frequency, general-purpose tactics and their specialized, high-specificity counterparts forms a robust strategy. The former ensures extensive exploration of the search space, while the latter precisely resolves the intricate logical dependencies central to establishing semantic equivalence.
@@ 5 Conclusion
In this work, we addressed the critical gap in evaluating autoformalized statements, where existing metrics fail to adequately balance semantic and structural information. We introduced ASSESS, a novel framework whose core contribution is the TransTED Similarity metric. By augmenting the structural comparison of operator trees with a set of semantic transformations, our method provides a more holistic and robust measure of statement similarity. We also developed EPLA, an expert-annotated benchmark for this task. Experiments conducted on EPLA demonstrate that TransTED Similarity establishes a new state-of-the-art, outperforming existing methods in both accuracy and Kappa score. This work provides the autoformalization community with a more reliable, efficient, and reproducible metric.
Acknowledgements
This work is sponsored by the National Key R&D Program of China Grant No. 2022YFA1008200 (T. L.). We also thank Shanghai Institute for Mathematics and Interdisciplinary Sciences (SIMIS) for their financial support. This research was funded by SIMIS under grant number SIMIS-ID-2025-ST. The authors are grateful for the resources and facilities provided by SIMIS, which were essential for the completion of this work. We appreciate the valuable assistance of Yifan Bai and Yantao Li with the human evaluation.
Reproducibility Statement
We are committed to ensuring the reproducibility of our research. To this end, we provide comprehensive details of our benchmark, experimental setup, and evaluation methodology. The code, benchmark and experimental results are available at https://github.com/XiaoyangLiu-sjtu/ASSESS.
Specific details for reproducing our results can be found in the following sections of the paper:
  • EPLA Benchmark: The construction methodology, data sources, generation process, and the complete expert annotation scheme for our benchmark are detailed in Section 4.1.
  • Baselines and Experiment Setting: The Lean toolchain version, implementation details and specific configurations for all baseline methods are provided in Section 4.2 and Appendix F.5.
  • Detailed Results: In addition to the main results in Table 3, detailed per-metric performance, including the number of True Positives, True Negatives, False Positives, and False Negatives, are available in Appendix F.2.
Appendix A The Use of Large Language Models (LLMs)
We utilized a large language model (LLM) as a general-purpose writing assistant during the preparation of this manuscript. The primary role of the LLM was to polish the text written by the authors, such as improving grammar. All intellectual contributions, including the research methodology, experimental results, and their interpretation, are entirely the work of the authors. The LLM’s function was strictly that of an assistive tool for improving the quality of the written presentation.
Appendix B Limitations and Future Work
We analyzed some failure cases for understanding TransTED Similarity metric’s boundaries and guiding future research. We categorize the primary failure modes into two types.
Limited Tactic Commands (False Negatives).
TransTED Similarity relies on a curated set of tactic commands (specifically rw? and 7 structural tactics). In complex cases requiring creative leaps or intermediate lemma synthesis, the heuristic search may fail to find a path, leading to a low score for equivalent statements.
EPLA-miniF2F #25
NL: Let $S=2010+2011+\cdots+4018$. Compute the residue of $S$, modulo 2009. Show that it is 0.
Label:
theorem mathd_numbertheory_353 (s : $\mathbb{N}$) (h0 : s = $\mathrm{\Sigma}$ k in Finset.Icc 2010 4018, k) : s % 2009 = 0 := by sorry
Prediction:
theorem my_favorite_theorem : (Finset.sum (Finset.Icc 2010 4018) id) % 2009 = 0 := by sorry
TED Similarity: $0.17$
TransTED Similarity: $0.17$
In this case, the current tactic set could not bridge the semantic gap between the variable definition s in the hypothesis and its direct usage in the conclusion. This highlights the need to further optimize the configuration of tactic commands.
Semantic-Structural Tension (False Positives).
Because TransTED Similarity includes a structural similarity component, it can assign high scores to statements that are structurally similar but semantically distinct (e.g., off-by-one errors or index shifts).
EPLA-ProofNet #360
NL: Prove that $\left(\sum_{j=1}^{n}a_{j}b_{j}\right)^{2}\leq\left(\sum_{j=1}^{n}ja_{j}{}^{2}\right)\left(\sum_{j=1}^{n}\frac{b_{j}{}^{2}}{j}\right)$ for all real numbers $a_{1},\ldots,a_{n}$ and $b_{1},\ldots,b_{n}$.
Label:
theorem exercise_6_3 {n : $\mathbb{N}$} (a b : Fin n $\rightarrow$ $\mathbb{R}$) : ($\mathrm{\Sigma}$ i, a i * b i) ^ 2 $\leq$ ($\mathrm{\Sigma}$ i : Fin n, i * a i ^ 2) * ($\mathrm{\Sigma}$ i, b i ^ 2 / i) := by sorry
Prediction:
theorem test_problem (n : $\mathbb{N}$) (a b : Fin n $\rightarrow$ $\mathbb{R}$) : ($\mathrm{\Sigma}$ j : Fin n, a j * b j)^2 $\leq$ ($\mathrm{\Sigma}$ j : Fin n, (j.val + 1) * (a j)^2) * ($\mathrm{\Sigma}$ j : Fin n, (b j)^2 / (j.val + 1)) := by sorry
TED Similarity: $0.51$
TransTED Similarity: $0.81$
The final transformed statement:
theorem node_4 (a : $\mathbb{N}$) (a_1 a_2 : Fin a $\rightarrow$ $\mathbb{R}$) :($\mathrm{\Sigma}$ i : Fin a, $\uparrow\uparrowi$ * a_1 i ^ 2) * $\mathrm{\Sigma}$ i : Fin a, a_2 i ^ 2 / $\uparrow\uparrowi$ = ($\mathrm{\Sigma}$ j : Fin a, ($\uparrow\uparrowj$ + 1) * a_1 j ^ 2) * $\mathrm{\Sigma}$ j : Fin a, a_2 j ^ 2 / ($\uparrow\uparrowj$ + 1) := by sorry
The label (left hand side) iterates over $i$, while the prediction (right hand side) iterates over $j+1$. Although TransTED Similarity fails to provide a correct binary verification on mathematical equivalence due to structural likeness, the searching result precisely locates the semantic difference. This suggests that TransTED Similarity might be highly effective for human-in-the-loop debugging, filtering nearly-correct candidates for expert review.
Furthermore, we concede that TransTED Similarity does not guarantee 100% precision. This is a fundamental characteristic of continuous metrics, which require a decision threshold and thus incur a precision-recall trade-off. Unlike rigorous decision metrics like BEq (which guarantee 100% precision but suffer from extremely low recall due to prover brittleness), TransTED Similarity is designed as a fractionalized metric. Its goal is to maximize alignment with human judgment by balancing precision and recall, rather than strictly filtering for absolute logical provability. Consequently, TransTED Similarity’s false positives are rarely meaningless. Instead, they are typically samples with “high semantic similarity but minor technical discrepancies,” such as:
  • Consistency issues in variable naming or type.
  • Differences in implicit type parameters.
  • The omission of a single minor hypothesis.
In practical scenarios like human-in-the-loop evaluation or dataset pre-screening, these similar but not correct candidates are highly valuable. A metric that flags these as highly similar helps human experts locate and correct minor errors, whereas a strict prover like BEq would simply reject them, providing no feedback.
Appendix C Proof of Theorem 1